# Guide-LLM: An Embodied LLM Agent and Text-Based Topological Map for Robotic Guidance of People with Visual Impairments

### v1

@@ IV Results
Localization Recovery 66%
IV-A Experiment Setup
We used iGibson [49] simulator to validate our framework. Two distinct environments of varying sizes were selected to evaluate the adaptability and robustness of our agent. These environments were chosen to represent different levels of complexity and scale, allowing us to test the agent’s performance in both small, confined spaces and larger, more open area. In each environment, random start and destination points represented in the text map were assigned for navigation. Turtlebot was equipped with RGB camera to capture images of current observation. All experiments were conducted on a laptop with an Intel Core I9 and 32 GB ram, and RTX 4070 GPU.
IV-B LLM Guided Navigation Ablation Study
The objective of this experiment was to assess whether our LLM agent could successfully navigate to a destination by interpreting user queries without detailed instructions, showcasing its potential to assist PVI. To evaluate the contributions of the core components in our agent framework, we performed an ablation study by systematically removing system prompts and path planning module individually to analyze the impact of each removal on navigation performance. As shown in Table I, when all components were active, Guide-LLM achieved an 83.33% success rate in navigation within an office environment. Removing the system prompt resulted in a 0% success rate, indicating that the system prompt is crucial for guiding the agent’s understanding and responses by providing necessary context and instructions.
To assess the impact of the path planning module, we removed it and directly input the text map into our agent. The success rate of navigation decreased to 40%. This decrease is attributed to several factors. First, directly providing the text map along with the system prompt increased the prompt length, which degraded the agent’s reasoning capability due to processing bottlenecks associated with large prompts [50, 51, 52]. Without path planning module, the agent is more successful when following alphabetical paths (e.g., Hallway A to Hallway C) but struggled with reversed paths (e.g., Hallway E to Hallway A). Removing the image retrieval system forced users to specify the starting location, altering the experiment and adding user burden. This configuration was excluded from comparative analysis to maintain experimental fairness, as it no longer tested the agent’s ability to infer starting points visually.
@@ IV-C Localization Error Detection and Recovery
To evaluate our agent’s error handling and recovery capabilities, we conducted an experiment where the robot was placed in random node during navigation. When the similarity score between the robot’s current observation and the expected navigational image fell below 0.94, the agent attempted to recover by querying the main vector database to re-localize itself and then restarted the navigation planning process from the recovered location. To further illustrate the system’s behavior, we provide an example chat box interaction in Fig. 3, showing how the agent detects the error, responds to the unexpected outcome, and initiates the recovery process by updating the plan based on the new location. This experiment was carried out 30 times in an office environment to thoroughly test the robustness of the agent. Experiment results are shown in Table.II To detect and recover from situations where the robot hits a wall, we modified the TurtleBot to output a message when it remains stationary for a certain amount of time after receiving a movement command. This allows the agent to recognize the issue and initiate a recovery process by adjusting the navigation plan. During the experiment, we observed that the localization error occasionally happens when the agent is in a visually similar area, but managed to recover from some cases using commonsense by considering the past path it took.
TABLE III: Hazard detection performance
Scenario Total detection True positive False positive
Hazard detection 30 12 18
@@ IV-D Commonsense Reasoning for Hazard Detection
The objective of this experiment is to evaluate the agent’s ability to identify potential hazards in the environment, communicate these risks to the user effectively, and adapt its actions based on the user’s decisions. To test this, images of obstacles and potential hazards were placed along a navigation path, including common dangers such as warning signs, physical barriers, and overhanging objects. The agent was evaluated on its ability to detect these hazards and provide appropriate warnings to the user, suggesting alternative routes when necessary. As shown in Table.III key metrics included hazard detection accuracy, with a focus on true positive rates and false positives, such as instances where the agent incorrectly identified non-hazardous objects (e.g., a can of coke near a wall) as potential dangers. An example scenario is reflected in Fig. 3 where the agent demonstrates its commonsense reasoning by going beyond predefined navigation instructions. Upon detecting potential hazards, the agent warns the user and provides possible alternatives. For example, if an obstacle blocks the path or if a risky surface (such as a wet floor sign) is identified, the agent alerts the user with a warning message and suggests a safer alternate route. Once the hazard is communicated, the system waits for the user’s response, allowing the user to either accept the suggested route or proceed with caution along the original path. This experiment highlights the agent’s capacity to dynamically reason and adapt to unforeseen environmental changes.
@@ IV-E Personalization Potential
Research on navigational aid preferences shows that personalization is crucial for improving the user experience [53]. The LLM-based system is capable of meeting these diverse preferences through natural language interaction, allowing users to easily communicate their specific needs. By simply talking to the agent, users can adjust parameters such as preferred walking speed, route type (e.g., avoiding stairs), or safety concerns (e.g., avoiding potential hazard). This flexibility in tailoring the experience highlights the personalization potential of the LLM system. To evaluate how user preferences are reflected, we re-named some spaces in text topological map for example, concert hall, food court, noisy area and quiet area. We evaluated the agent’s routing performance based on user preferences, testing it 10 times on the same route to determine if it made different choices. The results were as follows: users preferring detailed step-by-step navigation instructions during navigation had their preferences met 10 out of 10 times, those preferring quiet routes had their preferences met 10 out of 10 times, and users who preferred stairs over elevators saw their preferences met 3 out of 10 times. The primary reason for the high failure rate in the last experiment was the excessive number of route options available to reach the destination, which may have exceeded the maximum allowed prompt, leading to the failure. As shown in Fig.3, left corner of the map had rectangular shape causing path planning module produce 8 route options causing agent failing to reflect on user preferences.
@@ V Conclusion and Future Work
We introduced Guide-LLM, an innovative framework leveraging Large Language Models (LLMs) and novel text-based topological map to assist persons with visual impairments (PVI) in navigating large indoor environments. Our system successfully demonstrated the ability to provide efficient, adaptive, and personalized navigation, significantly reducing the need for detailed user instructions. Future work will focus on expanding the system’s capabilities, including autonomous exploration and map generation, as well as addressing real-time challenges such as obstacle avoidance. Testing in real-world scenarios with PVI will be on future works to improve and refine the system. These advancements will bring us closer to providing a comprehensive assistive solution that empowers PVI with greater independence and confidence in navigating complex environments.

### v2

@@ IV-B Model Selection and Topological Navigation
Recovery 72%
TABLE IV: Hazard Detection Results
Scenario # Trials TP FP FN TN
Potential Hazard 30 25 5
Non-Potential Hazard 30 10 20
@@ IV-C Localization Error Detection and Recovery
To evaluate Guide-LLM’s resilience against localization failures, we introduced the ‘kidnapped robot’ scenario in the office environment. During navigation, the robot was randomly teleported to a different node, forcing the model to detect error and recover from it. We performed this evaluation 50 times in total, ignoring naturally failed scenarios where navigation failed due to LLM’s navigation failure. If the agent successfully recognized the mismatch between its expected location and current observation, we considered the localization error detection successful. If the agent re-localize and reached the goal by updating its plan based on the new location, the trial was marked as a successful recovery. Table III shows the localization error detection and recovery rates. In addition, we observed that most failures occurred in feature-sparse corridors or visually similar places. We also observe that recovery failures happen when the Guide-LLM recognized the localization error but carries on with the previous path.
@@ IV-D Commonsense Reasoning for Hazard Detection
The objective of this experiment was to evaluate the agent’s ability to identify potential hazards in the environment, communicate these risks to the user, and adapt its actions based on user decisions. To test this capability, we placed obstacles and potential hazards (warning signs, physical barriers) and non-potential hazards along the navigation path (chair, pot). As shown in Table IV, we conducted 60 trials: 30 with objects and hazards with the potential to impede navigation, while the other 30 with non-hazardous objects that did not interfere with safe travel. In the hazard scenario, we recorded 25 true positives (TP), where the agent correctly identified a hazard and warned the user, and 5 false negatives (FN), where it failed to recognize a hazard and proceeded without warning. These false negatives typically resulted from limited height awareness of both the Guide-LLM and the user. In the non-hazard scenario, we observed 10 false positives (FP), where the agent mistakenly flagged hazard, and 20 true negatives (TN), where it ignored a harmless item. Figure 3 illustrates a typical hazard scenario in which the agent demonstrates commonsense reasoning beyond navigation. Upon detecting a hazard, it alerts the user and proposes an alternate route, underscoring the system’s capacity to dynamically adapt when faced with unexpected obstacles.
@@ IV-E Personalization Potential
Prior research on navigational aids indicates that personalization is vital for enhancing user experience [53]. As shown in Table I, current navigation aids are limited in their ability to incorporate personalized preferences. In contrast, Guide-LLM supports diverse user needs through natural language interaction, enabling users to easily communicate specific requirements. To illustrate how user preferences can be addressed, we renamed certain spaces in the text-based topological map (e.g., “concert hall,”, ”food court”, “noisy area,” “quiet area”). We then conducted 50 repeated trials on the same route, applying various user preferences and phrasings (e.g., avoiding stairs, noisy areas, or avoiding places likely to be crowded). In each trial, Guide-LLM adapted its path planning accordingly, even selecting routes up to three times longer if it aligned with the user’s stated requirements. This is possible by the integration of the text-based map with the path planning module, allowing the Guide-LLM to evaluate alternative routes and plan ahead. Some failures were seen when the number of possible routes was excessively large, potentially exceeding the maximum context window and leading to suboptimal decisions. This test resulted 92% success rate. Additionally, we evaluated Guide-LLM’s ability to adjust speed by providing multiple instructional cues (e.g., asking it to speed up or slow down) to see whether the system could infer user intent. We observed that Guide-LLM consistently increased or decreased walking speed without any failures. For characteristic changes such as providing detailed, step-by-step navigation or shifting conversational tone (e.g., formal, casual, or informal), the system met user expectations in 43 out of 50 trials.
IV-F LLM Guided Navigation Ablation Study
We conducted an ablation study to examine how each core component of the Guide-LLM framework contributes to navigation performance. As shown in Table II, the full system achieved an 86% success rate when navigating through an office environment. We then systematically removed individual components and measured the resulting performance drops. Removing the system prompt reduced the success rate to 0%, indicating that prompt-driven context is essential for the agent’s high-level understanding and instruction formulation. Next, we removed the path planning module and directly supplied the text-based map to the agent. This change lowered the success rate to 40%. The primary reason is that adding a large text map to the system prompt can exceed the context window, which degrades the agent’s reasoning capability due to processing bottlenecks associated with large prompts  [54, 55, 56]. Finally, removing the image retrieval system forced users to specify the starting location, thereby altering the experimental conditions and increasing user burden. Since it no longer tested the agent’s ability to visually infer starting points, we excluded this configuration from our comparative analysis to maintain fairness.
@@ V Conclusion and Future Work
We introduce Guide-LLM, an embodied large language model framework that can provide efficient, adaptive, and personalized navigation assistance for people with visual impairments. By combining a text-based topological map with common sense reasoning, our system reduces the need for explicit user instructions, enabling personalized navigation, offering a more intuitive navigational experience. Simulation results demonstrate that Guide-LLM excels in tasks such as localization error detection recovery, hazard detection, and personalization. As this research focused on developing and validating the technical feasibility of LLM based navigation aid, we plan to focus on the following directions in the future: Optimizing smaller language models to match or approximate the performance of larger ones, allowing on-device processing without reliance on cloud-based services. Expanding the text-based map creation process to autonomously generate and update environmental representations in large, complex buildings such as multi-story shopping center. Attaching depth camera to provide depth information to improve on hazard detection. Conducting real-world trials with visually impaired participants to evaluate usability, comfort, and overall effectiveness in navigation tasks. By addressing these challenges, we aim to refine Guide-LLM into a comprehensive assistive solution that empowers visually impaired individuals with more flexible, accessible, and safe navigation.